1. 1L sample of gas at 1atm pressure and 298K expands isothermally to 10L. It is then heated to 500K, compressed to 1L and then cooled 298K. What is the ∆U of overall the process?
1
Expert's answer
2020-07-09T14:56:01-0400

Solution.

The first process proceeds isothermically (at T=const), therefore, the change in internal energy is equal to 0. From the first process, we need to find out what pressure was formed by the SB using the klayperon-Mendeleev equation.

"p1 \times V1 = p2 \times V2"

p2 = 0.1 atm.

"p \times V = n \times R \times T"

"n = \frac{p \times V}{R \times T}"

n = 0.041 mol

The second process clearly occurs at constant pressure (p=const). Expression of the first law of thermodynamics for an Isobaric process:

"\Delta U = Q-A"

"A = -p \times \Delta V"

"A = -0.1*101325*(0.01-0.001) = -91.19 \ J"

"Q = \nu \times (\frac{5}{2}) \times R \times \Delta T"

"Q = 0.041 \times (\frac{5}{2}) \times 8.31 \times 202 = 172.06 \ J"

"\Delta U = 172.06 +91.19 = 263.25 \ J"

Answer:

"\Delta U = 263.25 \ J"